# Mind the Gap: Evaluating Vision Systems in Small Data Applications

### v1

@@ 1 Introduction
To systematically evaluate the small-data regime, we use the Natural World Tasks [NeWT; 38] benchmark, which is specifically designed for challenging fine-grained ecological classification tasks requiring expert annotation. Using this benchmark, we compare multimodal large language models (MLLMs) and vison-only methods across varying training set sizes. Using ecological classification tasks as representative test cases, we analyze model performance and scaling behavior. Our findings highlight significant limitations of current MLLMs, notably early performance plateaus, in contrast to sustained performance improvements observed in vision-only methods as dataset sizes increase within the small-data regime (see Fig. 2). While our study utilizes ecological tasks as a convenient testbed due to the availability of the NeWT benchmark in the small-data regime, our argued position extends beyond ecology to the broader field of computer vision applications where limited labeled data is a common constraint. In this work, we:
@@ 2 Background & Related Work
Recent computer vision methods [30, 28], and (multimodal) large language models [(M)LLMs; 27, 2, 11, 3] primarily evaluate performance using zero-shot or few-shot benchmarks [17, 34, 42]. These benchmarks emphasize generalization with extremely limited examples, reflecting a trend toward model robustness with minimal fine-tuning.
@@ 3 Methodology
To highlight the overlooked small-data regime in AI research, we introduce a rigorous experimental framework comparing MLLMs and vision-only models combined with traditional machine learning approaches specifically within this scenario. Unlike widely-studied zero-shot and few-shot benchmarks, our experiments explicitly target moderate data scales, ranging from tens to thousands of samples, to uncover non-obvious scaling behaviors and limitations of state-of-the-art methods. By providing detailed methodological guidance, we encourage researchers to adopt similar small-data evaluations, facilitating meaningful insights and practical recommendations for future method development.
@@
Specifically, we evaluate multiple state-of-the-art MLLMs (e.g., Gemini Flash, Qwen2.5-VL) alongside vision encoders (e.g., DINOv2, CLIP variants) paired with SVM-based classifiers on diverse ecological datasets from the NeWT benchmark. We systematically vary training set sizes and employs standardized prompting and parsing procedures to rigorously characterize model behaviors and scaling trends.
@@
We evaluate Gemini Flash 2.0, Gemini Flash 1.5 8B, Qwen2-VL 7B and Qwen2.5-VL 72B.
@@
We extract features from DINOv2, CLIP, and SigLIP. We test ViT-B, ViT-L, and ViT-H variants. The features are classified with an SVM; SVM hyperparameters are tuned using scikit-learn’s cross-validation grid search [29].
Tasks
We evaluate models on the NeWT benchmark [38]. NeWT contains ecologically-motivated binary classification tasks, each with to training samples. Tasks are grouped into eight clusters: species, attributes, health, ages, gestalt, contexts, counting and behavior. See both Appendix B and the original text for additional details.
Training Data
To analyze performance scaling, we train models on different amounts of labeled data. We define training subsets with sample sizes of , , , , , , , and all examples. Training examples are sampled uniformly and we ensure that there is at least one example per class when there are three or more examples.
@@
For all tasks, we compute bootstrapped confidence intervals by resampling test sets times with replacement and reporting the % confidence interval. MLLM responses are parsed using deterministic regex-based extraction. If multiple species are listed, we take the first species mentioned. Code to reproduce our results is available at https://github.com/samuelstevens/mindthegap.
@@ 4.1 Scaling Data
As shown in Fig. 3, MLLMs and vision-only methods exhibit fundamentally different scaling behaviors as training set size increases. MLLMs demonstrate rapid initial gains with very few examples (1-3) but consistently reach performance plateaus after 10-30 examples across most task clusters. In contrast, leveraging SVMs with vision transformers [ViTs; 12] show continuous, near-logarithmic improvement throughout the entire small-data regime, with no evidence of plateauing even at the maximum available training set size. This scaling disparity results in a widening performance gap as dataset size increases.
@@ 4.2 Scaling Models
Adding parameters to large language models demonstrates consistent improvement [20, 18, 41]. Our analysis reveals a different pattern for vision models in ecological tasks. As Fig. 4 illustrates, increasing computational resources yields diminishing returns compared to simply adding more training examples. Even as we scale SigLIP models across model and image sizes from to + GFLOPs, accuracy improvements remain modest, with a increase in training samples consistently outperforming a increase in computational capacity. This challenges the dominant “bigger is better” paradigm in recent AI research [24, 40].
@@ 4.3 Pre-Training Supervision
Our analysis reveals distinct performance patterns between vision-only and language-supervised pre-training approaches across ecological task clusters (Fig. 5). These differences underscore how pre-training objectives fundamentally influence a model’s capabilities.
@@
Conversely, language-supervised pre-training demonstrates substantial advantages in tasks requiring semantic understanding and contextual reasoning, most notably in ‘Gestalt’, ‘Context’, and ‘Behavior’ clusters, where models must recognize abstract visual concepts like image quality or complex animal activities. This pattern suggests that image-text learning provides semantic grounding that pure vision models struggle to develop.
@@ 5 Conclusion & Future Work
Our systematic evaluation of the small-data regime reveals distinct performance patterns: vision-only systems exhibit sustained improvement while MLLMs demonstrate early performance plateaus beyond 10-30 examples, suggesting that prompting struggles to learn nuanced patterns beyond a critical threshold of examples [19, 44]. Our findings underscore the critical importance of evaluating AI methods explicitly within the small-data regime, an evaluation practice largely overlooked in current work despite its relevance to real applications. By highlighting this evaluation gap, we hope to encourage more comprehensive benchmarking practices that better reflect the diverse data contexts encountered in practice.
@@
Appendix A: Methodology details for Sec. 3.
@@
Appendix B: NeWT details for Sec. 3.
@@
Appendix C: Additional results for Sec. 4.3.
@@ Appendix A Methodology Details
We test the following models, using API access where necessary: Gemini Flash 2.0, Gemini Flash 1.5 8B, Qwen2-VL 7B and Qwen2.5-VL 72B.
@@
Few-shot Sampling Methodology
Training subsets are generated with the following sample sizes: , , , , , , and where applicable. These subsets are sampled randomly from the full training split. Examples are sampled uniformly from the available training set.
Prediction Extraction from MLLMs
MLLM responses are parsed using deterministic regex-based extraction.
@@
Code to reproduce our results is available at https://github.com/samuelstevens/mindthegap.
@@ Appendix B Background on NeWT
Natural World Tasks [NeWT; 38] is a collection of binary classification tasks that go beyond species classification. The tasks are manually curated with a uniform distribution of positive and negative examples (so accuracy is an appropriate metric).

### v2

# Mind the (Data) Gap: Evaluating Vision Systems in Small Data Applications

@@ 1 Introduction
To systematically evaluate the small-data regime, we use the Natural World Tasks [NeWT; 42] benchmark, which is specifically designed for challenging fine-grained ecological classification tasks requiring expert annotation. Using this benchmark, we compare multimodal large language models (MLLMs, à la Qwen2.5-VL or Gemini Flash 2.0) and vision transformers (ViTs) across varying training set sizes. Using ecological classification tasks as representative test cases, we analyze model performance and scaling behavior. Our findings highlight significant limitations of current MLLMs, notably early performance plateaus, in contrast to sustained performance improvements observed in vision-only methods as dataset sizes increase within the small-data regime (see Fig. 1). While our study utilizes ecological tasks as a convenient testbed due to the availability of the NeWT benchmark in the small-data regime, our argued position extends beyond ecology to the broader field of computer vision applications where limited labeled data is a common constraint. In this work, we:
@@ 2 Background & Related Work
Recent computer vision methods [32, 30], and (multimodal) large language models [(M)LLMs; 29, 2, 11, 3] primarily evaluate performance using zero-shot or few-shot benchmarks [19, 38, 46]. These benchmarks emphasize generalization with extremely limited examples, reflecting a trend toward model robustness with minimal fine-tuning.
@@ 3 Methodology
To highlight the overlooked small-data regime in AI research, we introduce a rigorous experimental framework comparing multi-modal large language models (MLLMs) and vision transformers combined with traditional machine learning approaches specifically within this scenario. Unlike widely-studied zero-shot and few-shot benchmarks, our experiments explicitly target moderate data scales, ranging from tens to thousands of samples, to uncover non-obvious scaling behaviors and limitations of state-of-the-art methods. By providing detailed methodological guidance, we encourage researchers to adopt similar small-data evaluations, facilitating meaningful insights and practical recommendations for future method development.
@@
Specifically, we evaluate multiple state-of-the-art MLLMs (e.g., Gemini Flash, Qwen2.5-VL) alongside vision encoders (e.g., DINOv2, CLIP variants) paired with SVM-based classifiers on diverse ecological datasets from the NeWT benchmark. We systematically vary the number of labeled examples and apply standard prompting and parsing procedures to rigorously characterize model behaviors and scaling trends. Section 3 contain our prompts and pseudo-code for parsing responses.
Tasks
We evaluate models on the NeWT benchmark [42]. NeWT contains ecologically-motivated binary classification tasks, each with to labeled examples. Tasks are grouped into eight clusters: species, attributes, health, ages, gestalt, contexts, counting and behavior. See both Appendix B and the original text for additional details.
@@
We evaluate Gemini Flash 2.0, Gemini Flash 1.5 8B, Qwen2-VL 7B and Qwen2.5-VL 72B. At the time of writing, these models represent the best tradeoff between cost and performance on a variety of benchmarks.
@@
We extract features from DINOv2, CLIP, and SigLIP, three popular pre-trained, general-domain vision encoders. We test ViT-B, ViT-L, and ViT-H variants. Per NeWT’s original methodology, we exclusively use SVMs as our binary classifier on top of dense vision model features; SVM hyperparameters are tuned using scikit-learn’s cross-validation grid search [31].
Labeled Examples
To analyze performance scaling, we train models on different amounts of labeled examples. We define near-logarithmically scaled training subsets with sample sizes of , , , , , , , and all examples. Labeled examples are sampled uniformly and we ensure that there is at least one example per class when there are two or more examples.
@@
For all tasks, we compute bootstrapped confidence intervals by resampling test sets times with replacement and reporting the % confidence interval. MLLM responses are parsed using deterministic regex-based extraction. If multiple species are listed, we take the first species mentioned.
@@ 4.1 Scaling Data
As shown in Fig. 2, MLLMs and vision-only methods exhibit fundamentally different scaling behaviors as the number of labeled examples increases. MLLMs demonstrate rapid initial gains with very few examples (1-3) but consistently reach performance plateaus after 10-30 examples across most task clusters. In contrast, leveraging SVMs with vision transformers [ViTs; 13] show continuous, near-logarithmic improvement throughout the entire small-data regime, with no evidence of plateauing. This scaling disparity results in a widening performance gap as dataset size increases.
@@ 4.2 Scaling Models
Adding parameters to large language models demonstrates consistent improvement [22, 20, 45]. Our analysis reveals a different pattern for vision models in ecological tasks. As Fig. 3 illustrates, increasing computational resources yields diminishing returns compared to simply adding more labeled examples. Even as we scale SigLIP models across model and image sizes from to + GFLOPs, accuracy improvements remain modest, with a increase in labeled samples consistently outperforming a increase in computational capacity. This challenges the dominant “bigger is better” paradigm in recent AI research [26, 44].
Several factors could explain this difference from language model scaling properties. The emergence threshold for vision models might occur at parameter counts beyond our experimental range [36, 15, 37, 12]. Pretraining methodology differences are significant—vision models employ diverse objectives (contrastive, self-supervised, supervised) compared to the converged next-token prediction approach in language. Our findings indicate that for ecological computer vision tasks within the small-data regime, prioritizing data collection provides more reliable performance improvements than scaling computational resources alone.
@@ 4.3 Pre-Training Supervision
We use these results to reveal distinct performance patterns between vision-only and language-supervised pre-training approaches across ecological task clusters (Fig. 3). These differences underscore how pre-training objectives fundamentally influence a model’s capabilities.
@@
Conversely, language-supervised pre-training demonstrates substantial advantages in tasks requiring semantic understanding and contextual reasoning, notably in ‘Gestalt’, ‘Context’, and ‘Behavior’ clusters, where models must recognize abstract visual concepts like image quality or animal activities. This suggests that image-text learning provides semantic grounding that pure vision models lack.
@@ 5 Conclusion & Future Work
Our systematic evaluation of the small-data regime reveals distinct performance patterns: vision-only systems exhibit sustained improvement while MLLMs demonstrate early performance plateaus beyond 10-30 labeled examples, suggesting that prompting struggles to learn nuanced patterns beyond a critical threshold of examples [21, 48]. Our findings underscore the critical importance of evaluating AI methods explicitly within the small-data regime, an evaluation practice largely overlooked in current work despite its relevance to real applications. By highlighting this evaluation gap, we hope to encourage more comprehensive benchmarking practices that better reflect the diverse data contexts encountered in practice.
@@
Appendix A: Methodology details for Section 3.
@@
Appendix B: NeWT details for Section 3.
@@
Appendix C: Additional results for Section 4.3.
@@ Appendix A Methodology Details
We test the following models via API access: Gemini Flash 2.0, Gemini Flash 1.5 8B, Qwen2-VL 7B and Qwen2.5-VL 72B.
@@
Sampling:
Subsets of labeled examples are sampled uniformly from the full training split with the following sample sizes: , , , , , , and where applicable.
Prompting and Parsing:
All the tasks in NeWT are binary classification tasks. We use the following template: “What is this a picture of, ’{a}’ or ’{b}’? Respond with your answer in bold.” where {a} and {b} are replaced with the two classes, with a random order. MLLM responses are parsed using regex-based extraction, using character-based distance to pick the classname closest to whatever bold text is first found in the response.
@@ Appendix B Background on NeWT
Natural World Tasks [NeWT; 42] is a collection of binary classification tasks that go beyond species classification. The tasks are manually curated with a uniform distribution of positive and negative examples (so accuracy is an appropriate metric).
@@ Appendix D All Results
We include all results in Tables 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17 and 18. These results will also be made available in a machine-readable format.